You have space in your garden for 55 small flowering bulbs. Crocus bulbs cost $0.35 each and daffodil bulbs cost $0.65 each. Your budget allows you to spend $29.15 on bulbs. How many crocus bulbs and daffodil bulbs can you buy?
Answer by mananth(16949)

x Crocus numbers
y Daffodils numbers

x+y =55 .............1
0.35 x+0.65 y = 29.15 .............2
Eliminate y
multiply (1)by -0.65
Multiply (2) by 1
-0.65 x -0.65 y = -35.75
0.35 x + 0.65 y = 29.15
Add the two equations
-0.30 x = -6.60
/ -0.30
x = 22
plug value of x in (1)
x+y = 55.00
22 +y = 55
y = 55-22
y = 33

22 Crocus
33 Daffodils
m.ananth@hotmail.ca